Help I cant figure this out!!! Please break it down fo rme so i can understand it. Thank you
Solve the system of equations using the substitution method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
-7x + y = 4
2x + 8y = 0

Answer by stanbon(75887)

-7x + y = 4
2x + 8y = 0
-------------
Solve the 1st equation for Y;
y=7x+4
Substitute that into the 2nd equation to solve for x:
2x+8(7x+4)=0
2x+56x+32=0
58x = -32
x=-16/29
--------
Substitute that in y=7x+4 to solve for y:
7(-16/29)+4=-112/29 + 116/29 = 4/29
---------
Solution:
x=-16/29 ; y=4/29
============
Cheers,
Stan H.